OpenStudy (anonymous):

Trig/Pre Cal/identities Where do I start with this? Do not give answer. Just need to know where to start? I am not sure. \( \frac{cos x}{1-sin x}=\frac{1+sin x}{cos x} \) Thank you

OpenStudy (anonymous):

I would multiply the left side by \[\frac{ 1 + \sin x }{ 1 + \sin x}\] because \[\cos^2 x = 1 - \sin^2 x\]

OpenStudy (anonymous):

\( 1- \cot x=\frac{1+\sin x}{\cos x} \) ?

OpenStudy (anonymous):

Ok, let me work it from there

OpenStudy (anonymous):

Yep that is it. Thank you!!!!!

OpenStudy (anonymous):

@peachpi I sent ya a owlbuck too ;-) thanks

OpenStudy (anonymous):

you're welcome. and thank you for the owlbuck!